### v1

@@ 3.5 Acceptance Tests by JUNO
A warehouse designated for the JUNO 20-inch PMT mass testing [5] was also used for the storage and acceptance tests of the 3-inch PMT cables and connectors. The initial step involved visual inspections to detect any visible damage or defects, with particular attention paid to scratches on the sealing surfaces of the connectors. Electrical tests followed, assessing the resistance and electrical conduction of the core wire and the braid grounding wire, as well as verifying the channel sequence from 1 to 16. Additionally, HV testing (1.6 kV, 5 minutes) was carried out to determine the cables’ ability to withstand voltage stress and confirm their safe operation under normal conditions. Since only two receptacles from each batch underwent destructive tests underwater at Axon’, a supplementary leak test with SF6 gas was conducted for 10% of the receptacles. Each selected sample was sealed with a fake plug using an axial O-ring, evacuated to below 50 Pa, maintained for one minute, and then filled with SF6 gas at a pressure about 25% higher than standard atmospheric pressure. An SF6 detector was used outside the receptacle to detect any leaked gas with a sensitivity of 10Pam3/s [30]. No leakage was detected in any of the tested receptacles.
@@ 4.1 PMT Grouping
PMTs are made of quartz glass with an average thickness between 1 mm and 2 mm. During long-term operation in water at a 42 m depth, buckling failure may happen due to large external pressure. For an ideal spherical shell, the classical critical buckling pressure can be expressed as [31]
where is the Poisson’s ratio, is the tensile modulus, and
and are the radius and thickness of the glass shell, respectively. Since and are constants for the PMT glass shell, and the shape () is almost fixed, the critical pressure is sensitive to the thickness of the glass and thus to the weight.
A dedicated study was done to find the dependence of the critical pressure on the weight of the PMT. Ten PMTs were randomly selected from an early production batch and their weights were measured. All pins on the bottom of the PMT were sealed with polyurethane and each PMT was individually put into a water tank. The water pressure was increased with a water pump at a speed of about 0.5 MPa/min. When the PMT glass imploded, the water pressure was recorded.
Critical pressures at different PMT weights are shown in the bottom panel of Fig. 8. A positive correlation can be seen. However, there is a large spread from 1.5 MPa to 2.5 MPa in the 110 g to 115 g range, which can be attributed to variations in the glass shell thickness, as implosions typically initiate at points with the smallest thickness or highest stress.
A parabolic function (red dashed curve) is used to fit the critical pressure distribution, but with large fluctuation due to the special shape (not ideal spherical) and non-uniform thickess of the glass shell, as well as the weight of electric components inside the shell.
Figure 8: The top panel shows the measured weight distribution for all 26,000 PMTs. W1, W2, and W3 represent three groups of PMTs according to their weights. The blue points in the bottom panel are the measured critical pressures underwater as a function of the PMT weight. The red dashed line is the best fit parabola.
@@
They were categorized into three classes based on their weights and installed at maximum water depths of 10 m, 25 m, and 42 m, respectively, as detailed in Tab. 2. The critical pressure for each class is approximately five times greater than the expected water pressure. Only 3.8% of PMTs fall into class W1 and were installed near the top of the JUNO detector at less than 10 m depth. About 40.5% PMTs, belonging to class W2, were installed on the upper hemisphere of the JUNO detector at depths smaller than 25 m. The remaining PMTs, comprising more than half of the total, belong to class W3 and were installed mostly on the lower hemisphere.
Table 2: Classification of PMTs by weight.
@@ 4.2 Waterproof PMT Potting
The schematic design of the waterproof PMT potting is illustrated in Fig. 9. The PMT pins, the HV divider, and the cable are soldered and enclosed in an ABS plastic shell. This shell is filled with polyurethane to protect the electrical components from water and to block any light that might be emitted spontaneously by electrical discharges. The seal between the shell and the PMT glass is achieved using butyl tape, which is encompassed by a shrinkable tube and finished with low-pressure injection molding at the end.
@@
As a final verification of the waterproof potting design, 50 samples were produced using PMT glass bulbs or discarded PMTs along with HV dividers and short cables. The cable ends were sealed with epoxy, and all samples were submerged in a water tank in four separate batches, subjected to a pressure of 0.6 MPa for a minimum of 7 days. A potential difference of 2,000 V was applied through the cable ends, and currents were measured both before and after the high-pressure water test. The observed differences ranged from -40% to 70%, which can be attributed to variations in PMT exposure to light. If water had penetrated into the PMT pins or the divider, the current would have increased significantly, by at least an order of magnitude.
The primary steps involved in the potting are illustrated in Fig. 10 (a). An injection-molded ABS plastic shell was designed to provide mechanical support. Initially, the cable was threaded through a bottom entrance with a diameter of 2.2 mm, ensuring 28 cm of the cable remained inside the shell. Following this, epoxy was injected from the top to offer a bonding length greater than 8 cm, serving as the inner seal. The epoxy used is specialized for HDPE, featuring a bonding strength of 1.6 MPa. The sealing was further reinforced outside the shell’s tail through low-pressure injection molding with a polyamide and polyolefin mixture, also customized for HDPE.
Figure 10: (a) Major steps of waterproof PMT potting: 1, cable sealing with epoxy; 2, cable sealing with injection molding; 3, soldering HV divider on the bare PMT; 4, assembly of butyl tape around the PMT; 5, filling polyurethane; 6, inverting the PMT; 7, covering the shrinkable tube. The PMT signals were verified after step 3 and step 7 with HV. (b) A group of 16 PMTs with cables and the connector after waterproof potting.
Subsequently, the PMT was mounted and soldered onto the HV divider along with the cable. A set of 16 PMTs with plugs were then transferred to a dark box, where each channel was connected to a splitter box to separate the HV and the signal. A nominal HV was applied to each channel, and the current was monitored using an ammeter. Utilizing low-intensity LED lights inside the dark box, PMT signals, predominantly SPEs, were captured by an oscilloscope. The amplitude was required to be within 50% of the average. If a significant deviation was detected, the PMT was replaced, and the process was repeated for a new PMT.
@@ 4.3 Leakage Test
A large steel water tank was utilized to conduct a sampling test on potted PMTs to verify water leakage, as depicted in Fig. 11. A group of 16 PMTs was secured on a rack, with the plug sealed using two O-rings within a customized receptacle. Nine groups of PMTs of identical weight classification were simultaneously placed in the tank. The water pressure was elevated according to the weight specifications listed in Table 2 and maintained for 48 hours. Subsequently, the PMTs were removed from the tank, rinsed with pure water, and dried in an oven at 55oC. The PMTs were then subjected to HV testing, and signals were measured for the third time in the dark box.
@@ 5.2 Acceptance Criteria
The characterization of bare PMTs was conducted at the factory of Hainan Zhanchuang Photonics Technology Co., Ltd (HZC) [32], as noted in Ref. [6].
@@ 5.3 HV Calibration Testing
This value was then compared with the working HV determined from bare tube testing by the PMT manufacturer HZC (referred to HZC HV, with each increment of 25 V represents a distinct HV level). This comparison resulted in a linear relationship, as depicted in Fig. 13.
@@ 5.4 Test Results
Another important acceptance parameter is the DCR, which is measured by the threshold scanning method. In this test, the number of trigger signals from each PMT channel in an electronics time period of 0.419 s is measured as a function of the trigger threshold in DAC units ranging from 0 to 870. The trigger values are in inverse order: a high DACu value corresponds to a low trigger threshold value in amplitude. The counts versus the discriminator threshold (also called S-curve) is shown in Fig. 18. The PMT DCR was obtained by calculating the trigger rate at threshold 870 DACu, which corresponds to 0.3 PE. Figure 19 shows the DCR distribution of the potted PMTs and the bare PMTs. The average DCR for the bare PMTs is approximately 510 Hz, with a significant tail in the distribution primarily attributed to insufficient time spent in the dark room under high voltage (HV) for the DCR to stabilize before testing. For instance, some PMTs were only subjected to a 2-hour HV load, and a DCR value below 3 kHz was considered acceptable. The potted PMTs benefited from sufficient cooling time with working HV of 8 hours, resulting in a significant reduction in the tail component and an average dark count rate of approximately 420 Hz, lower than that observed for the bare PMT test in HZC. However, upon comparing the main peak position of the dark count rate distribution, a significant shift is observed in the acceptance test, with the peak position approximately 100 Hz higher than that in the bare PMT test. Our investigation reveals that this additional contribution primarily comes from fluorescence emitted by the packaging foam employed to protect the PMTs. Subsequent tests conducted on small batches of PMTs demonstrated a difference of around 100 Hz in dark count rates between foam-protected and non-foam conditions. Given that the primary objective of the acceptance test is to validate the functionality of the potted PMTs, it was deemed unnecessary to conduct extensive testing without foam protection in order to ensure efficiency and prevent any damage during this process.
@@
Table 4: The result of acceptance test.
@@ 6 Summary
The instrumentation of the 3-inch PMTs of JUNO underwent rigorous testing at every stage of development and mass production. Each PMT was equipped with a divider capable of supplying up to 1,300 V across the various dynodes. Capacitors with wires and surface-mounted resistors were meticulously chosen to meet space constraints while ensuring reliability. A standard RG178 coaxial cable, available in 5 m or 10 m lengths, was used to transmit both HV and the signal. The cable jacket was customized to enhance waterproofing. A 16-channel waterproof connector, developed in collaboration with the Axon’ cable company, served as the interface between the PMTs and the electronics. Signal attenuation in the cable was adequately compensated by the voltage, and the observed analog crosstalk was below 0.3%. Two long-term underwater validations were conducted, demonstrating promising reliability with FIT values of less than 500 and 350, respectively, in line with JUNO’s requirements.

### v2

@@ 3.5 Acceptance Tests by JUNO
A warehouse designated for the JUNO 20-inch PMT mass testing [1] was also used for the storage and acceptance tests of the 3-inch PMT cables and connectors. The initial step involved visual inspections to detect any visible damage or defects, with particular attention paid to scratches on the sealing surfaces of the connectors. Electrical tests followed, assessing the resistance and electrical conduction of the core wire and the braid grounding wire, as well as verifying the channel sequence from 1 to 16. Additionally, HV testing (1.6 kV, 5 minutes) was carried out to determine the cables’ ability to withstand voltage stress and confirm their safe operation under normal conditions. Since only two receptacles from each batch underwent underwater tests at Axon’, a supplementary leak test with SF6 gas was conducted for 10% of the receptacles. Each selected sample was sealed with a fake plug using an axial O-ring, evacuated to below 50 Pa, maintained for one minute, and then filled with SF6 gas at a pressure about 25% higher than standard atmospheric pressure. An SF6 detector was used outside the receptacle to detect any leaked gas with a sensitivity of 10Pam3/s [16]. No leakage was detected in any of the tested receptacles.
@@ 4.1 PMT Grouping
PMTs are made of quartz glass with an average thickness between 1 mm and 2 mm. During long-term operation in water at a 42 m depth, buckling failure may happen due to large external pressure. For an ideal spherical shell, the classical critical buckling pressure in MPa can be expressed as [31]
where is the Poisson’s ratio, a dimensionless quantity, is the tensile modulus expressed in MPa, and
and are the radius and thickness of the glass shell in centimeter, respectively. Since and are constants for the PMT glass shell, and the shape (ratius ) is essentially fixed, the critical pressure is sensitive to the thickness of the glass and thus to the weight.
A dedicated study was made to find the dependence of the critical pressure on the weight of the PMT. Ten PMTs were randomly selected from an early production batch and their weights were measured. All pins on the bottom of the PMT were sealed with polyurethane and each PMT was individually put into a water tank. The water pressure was increased with a water pump at a speed of about 0.5 MPa/min. When the PMT glass imploded, the water pressure was recorded.
Critical pressures in comparison to atmospheric pressure at varying weights of PMTs are presented in the lower panel of Fig. 8. A positive relationship is evident. However, there is a large spread from 1.5 MPa to 2.5 MPa in the 110 g to 115 g range, which can be attributed to variations in the glass shell thickness, as implosions typically initiate at points with the smallest thickness or highest stress.
The obtained data was fitted with either a parabolic function as denoted in Eq. 1 or a linear function, giving comparable fitting qualities. A critical pressure exceeding 1 MPa was observed for glass bulbs weighing over 105 g when extrapolating using either of the two fitting functions.
Figure 8: The top panel shows the measured weight distribution for all 26,000 PMTs. W1, W2, and W3 represent three groups of PMTs according to their weights. The blue points in the bottom panel are the measured critical pressures underwater as a function of the PMT weight.
@@
They were categorized into three classes based on their weights and installed at maximum water depths of 10 m, 25 m, and 42 m, respectively, as detailed in Tab. 2. Only 3.8% of PMTs fall into class W1 and were installed near the top of the JUNO detector at less than 10 m depth. About 40.5% PMTs, belonging to class W2, were installed on the upper hemisphere of the JUNO detector at depths smaller than 25 m. The remaining PMTs, comprising more than half of the total, belong to class W3 and were installed mostly on the lower hemisphere. The critical pressure for W2 and W3 is three times greater than the anticipated water pressure. Although the safety margin is lower for W1, the proportion of PMTs in this category is minimal.
Table 2: Classification of PMTs based on weight. The water pressure indicated in the final column is presented relative to atmospheric pressure.
@@ 4.2 Waterproof PMT Potting
The schematic design of the waterproof PMT potting is illustrated in Fig. 9. The PMT pins, the HV divider, and the cable are soldered and enclosed in an Acrylonitrile Butadiene Styrene (ABS) plastic shell. This shell is filled with polyurethane to protect the electrical components from water and to block any light that might be emitted spontaneously by electrical discharges. The seal between the shell and the PMT glass is achieved using butyl tape, which is encompassed by a shrinkable tube and finished with low-pressure injection molding at the end.
@@
As a final validation of the waterproof potting design, 50 samples were fabricated using PMT glass bulbs or discarded PMTs, in combination with HV dividers and short cables. The cable terminations were sealed with epoxy, and the samples were immersed in a water tank in four separate batches, and subjected to a pressure of 0.6 MPa for a minimum of seven days. A voltage difference of 2,000 V was applied across the cable terminations by cutting the cable ends and removing the epoxy seal. The currents were recorded both before and after the high-pressure water test. For the glass bulb samples, the current measurements remained consistent. For the still functional discarded PMT samples, variations in current levels were noted, ranging from -40% to 70%, which could be attributed to differences in PMT light exposure. If water had infiltrated the PMT pins or the divider, a substantial increase in current, at least by an order of magnitude, would have been detected. Consequently, no instances of water leakage were detected in any of the tested samples.
The primary steps involved in the potting are illustrated in Fig. 10 (a). An injection-molded ABS plastic potting shell was designed to provide mechanical support. Initially, the cable was threaded through a bottom entrance with a diameter of 2.2 mm, ensuring 28 cm of the cable remained inside the shell. Following this, epoxy Loctite ES4412 [22] was injected from the top to offer a bonding length greater than 8 cm, serving as the inner seal. The epoxy used is specialized for HDPE, featuring a bonding strength of 1.6 MPa. The sealing was further reinforced outside the shell’s tail through low-pressure injection molding with a polyamide and polyolefin mixture, also customized for HDPE.
Figure 10: (a) Major steps of waterproof PMT potting: 1, cable sealing with epoxy; 2, cable sealing with injection molding; 3, soldering HV divider on the bare PMT; 4, assembly of butyl tape around the PMT; 5, filling polyurethane; 6, inverting the PMT; 7, covering with shrinkable tube. The PMT signals were verified after step 3 and step 7 with HV. (b) A group of 16 PMTs with cables and the connector after waterproof potting.
Subsequently, the PMT was mounted and soldered onto the HV divider along with the cable. A set of 16 PMTs with plugs were then transferred to a dark box, where each channel was connected to a splitter box to separate the HV and the signal. A nominal HV was applied to each channel, and the current was monitored using a micro-ammeter. Utilizing low-intensity LED lights inside the dark box, PMT signals, predominantly SPEs, were captured by an oscilloscope. The amplitude was required to be within 50% of the average. If a significant deviation was detected, the PMT was replaced, and the process was repeated for a new PMT.
@@ 4.3 Leakage Test
A large steel water tank was utilized to conduct a sampling test on potted PMTs to verify water leakage, as depicted in Fig. 11. A group of 16 PMTs was secured on a rack, with the plug sealed using two O-rings within a customized receptacle. Nine groups of PMTs of identical weight classification were simultaneously placed in the tank. The relative water pressure was elevated according to the weight specifications listed in Table 2 and maintained for 48 hours. Subsequently, the PMTs were removed from the tank, rinsed with pure water, and dried in an oven at 55oC. The PMTs were then subjected to HV testing, and signals were measured for the third time in the dark box.
@@ 5.2 Acceptance Criteria
As stated in Sec. 1, the bare PMTs were characterized in HZC.
@@ 5.3 HV Calibration Testing
This value was then compared with the working HV determined from bare tube testing by the PMT manufacturer HZC (referred to HZC HV, with each increment of 25 V representing a distinct HV level). This comparison resulted in a linear relationship, as depicted in Fig. 13.
@@ 5.4 Test Results
Another important acceptance parameter is the DCR, which is measured by the threshold scanning method. In this test, the number of trigger signals from each PMT channel in an electronics time period of 0.419 s is measured as a function of the trigger threshold in DAC units ranging from 0 to 870. The trigger values are in inverse order: a high DACu value corresponds to a low trigger threshold value in amplitude. The counts versus the discriminator threshold (also called the S-curve) is shown in Fig. 18. The PMT DCR was obtained by calculating the trigger rate at threshold 870 DACu, which corresponds to 0.3 PE. Figure 19 shows the DCR distribution of the potted PMTs and the bare PMTs. The average DCR for the bare PMTs is approximately 510 Hz, with a significant tail in the distribution primarily attributed to insufficient time spent in the dark room under high voltage (HV) for the DCR to stabilize before testing. For instance, some PMTs were only subjected to a 2-hour HV load, and a DCR value below 3 kHz was considered acceptable. The potted PMTs benefited from sufficient cooling time with working HV of 8 hours, resulting in a significant reduction in the tail component and an average dark count rate of approximately 420 Hz, lower than that observed for the bare PMT test in HZC. However, upon comparing the main peak position of the dark count rate distribution, a significant shift is observed in the acceptance test, with the peak position approximately 100 Hz higher than that in the bare PMT test. Our investigation reveals that this additional contribution primarily comes from fluorescence emitted by the packaging foam employed to protect the PMTs. Subsequent tests conducted on small batches of PMTs demonstrated a difference of around 100 Hz in dark count rates between foam-protected and non-foam conditions. Given that the primary objective of the acceptance test is to validate the functionality of the potted PMTs, it was deemed unnecessary to conduct extensive testing without foam protection in order to ensure efficiency and prevent any damage during this process.
@@
Table 4: Results of the acceptance tests for specific criteria, based on a total of 25,744 tested PMTs.
@@ 6 Summary
The instrumentation of the 3-inch PMTs of JUNO underwent rigorous testing at every stage of development and mass production. Each PMT was equipped with a divider capable of supplying up to 1,300 V across the various dynodes. Capacitors with wires and surface-mounted resistors were meticulously chosen to meet space constraints while ensuring reliability. A standard RG178 coaxial cable, available in 5 m or 10 m lengths, was used to transmit both HV and the signal. The cable jacket was customized to enhance waterproofing. A 16-channel waterproof connector, developed in collaboration with the Axon’ cable company, served as the interface between the PMTs and the electronics. Signal attenuation in the cable was adequately compensated by the voltage, and the observed analog crosstalk was below 0.3%. Two long-term underwater validations were conducted, demonstrating promising reliability with FIT values of less than 500 and 300, respectively, in line with JUNO’s requirements.